# Topological constraints on clean Lagrangian intersections from $\mathbb{Q}$-valued augmentations

### v2

@@ Abstract
Let $K$ be the connected sum in $\mathbb{R}^{3}$ of an arbitrary knot and the $(2,q)$-torus knot for $q\neq\pm 1$, and consider its conormal bundle $L_{K}$ in $T^{*}\mathbb{R}^{3}$. We prove that there is no compactly supported Hamiltonian diffeomorphism $\varphi$ on $T^{*}\mathbb{R}^{3}$ such that $\varphi(L_{K})$ and the zero section $\mathbb{R}^{3}$ intersect cleanly along the unknot in $\mathbb{R}^{3}$. For the proof, we examine the knot DGA, whose homology was proved to be isomorphic to the knot contact homology by Ekholm, Etnyre, Ng and Sullivan, with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$. We find an algebraic constraint imposed on every $\mathbb{Q}$-valued augmentation of the knot DGA of $K$ when $q\geq 3$. This constraint involves rational roots of polynomials in $\mathbb{Q}[T]$, and we apply Hilbert’s irreducibility theorem to complete the proof. Using a similar idea, we also prove an analogous result replacing the $(2,q)$-torus knot with the figure-eight knot.
@@ 1 Introduction
Background. Let us consider the cotangent bundle $T^{*}\mathbb{R}^{3}$ of $\mathbb{R}^{3}$ with a canonical symplectic structure. We identify the image of the zero section with $\mathbb{R}^{3}$. Given any knot $K$ in $\mathbb{R}^{3}$, its conormal bundle is defined by
$L_{K}\coloneqq\{(q,p)\in T^{*}\mathbb{R}^{3}\mid q\in K,\ \langle p,v\rangle=0\text{ for every }v\in T_{q}K\}.$
It is a Lagrangian submanifold of $T^{*}\mathbb{R}^{3}$ and intersects $\mathbb{R}^{3}$ cleanly along the knot $K$. Let $\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$ denote the set of Hamiltonian diffeomorphisms on $T^{*}\mathbb{R}^{3}$ associated to compactly supported Hamiltonians on $T^{*}\mathbb{R}^{3}$. Using these notations, we ask the following question, which was observed by the author in [16].
Question 1.1 (Question 1.1 of [16]).
Is there $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K})$ and $\mathbb{R}^{3}$ have a clean intersection along a knot in $\mathbb{R}^{3}$ whose isotopy class is different from $K$?
We remark that it is essential that $\varphi$ is a Hamiltonian diffeomorphism. Indeed, for any pair of knots $(K_{0},K_{1})$ in $\mathbb{R}^{3}$, there exists a $C^{\infty}$ isotopy $(\varphi^{t})_{t\in[0,1]}$ on $T^{*}\mathbb{R}^{3}$ with compact support such that $\varphi^{1}(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along $K_{1}$. See [16, Proposition 4.1].
As a special case, the answer to Question 1.1 when $K$ is the unknot is ‘no’. This can be deduced from the result by Ganatra-Pomerleano [9, Proposition 6.29] about clean Lagrangian intersections in a plumbing of two copies of the cotangent bundle of $S^{3}$ along the unknots (see [16, Section 1]). Smith-Wemyss [20, Proposition 4.10] and Asplund-Li [1, Theorem 1.3] also studies clean Lagrangian intersections in plumbings of two copies of $T^{*}S^{3}$ along the unknots. These results are concerned with Smith’s question on the ‘persistence (or rigidity) of unknottedness of Lagrangian intersections’ [9, Question 1.5].
In a general case, under the assumption that $\varphi(L_{K})$ for $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ intersects $\mathbb{R}^{3}$ cleanly along a knot $K^{\prime}$ in $\mathbb{R}^{3}$, the author proved [16, Theorem 1.1]. It relates the augmentation variety in $(\mathbb{C}^{*})^{2}$ of $K$ with that of $K^{\prime}$. This result leads us to the same answer to Question 1.1 as above when $K$ is the unknot [16, Theorem 1.3]. We can also deduce some constraints on the knot type of $K^{\prime}$ when $K$ has certain non-trivial knot types [16, Theorem 1.4, 1.5].
Main results. In fact, the aforementioned result [16, Theorem 1.1] is useless to solve the following weaker version of Question 1.1. For the reason, see Remark 1.5 below.
Question 1.2.
@@
The goal of this paper is to give a negative answer to this question when $K$ has a certain knot type. As a notation, let $K_{1}\#K_{2}$ denote the connected sum of knots $K_{1}$ and $K_{2}$ in $\mathbb{R}^{3}$. Let $T_{(2,2m+1)}$ for $m\in\mathbb{Z}$ denote the $(2,2m+1)$-torus knot and let $4_{1}$ denote the figure-eight knot.
@@
For the proof, the geometric arguments in this paper are similar to those in [16], which are outlined as follows: Let $K_{0}$ and $K_{1}$ be knots in $\mathbb{R}^{3}$. For $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$, suppose that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ intersect cleanly along $K_{1}$. We use the results by Ekholm-Etnyre-Ng-Sullivan [6, 3] which show that the Chekanov-Eliashberg DGA of the unit conormal bundle of $K_{i}$ is stable tame isomorphic to the knot DGA of $K_{i}$ for $i=0,1$. The knot DGA, introduced by Ng [12], is defined combinatorially from a braid representation of a knot. Here, we refer to the one with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$ defined in [14, 3, 15]. We then apply a DGA map associated to an exact Lagrangian cobordism between the unit conormal bundles of $K_{0}$ and $K_{1}$. This is a part of the symplectic field theory introduced by Eliashberg-Givental-Hofer [8]. Consequently, we obtain a DGA map from the knot DGA of $K_{0}$ to that of $K_{1}$. From a simple description of the knot DGA of the unknot, we obtain the following result about augmentations of the knot DGA of $K_{0}$.
Proposition 1.2 (Proposition 3.5).
Let $\mathbb{k}$ be a field. Suppose that $K_{1}$ is the unknot. Then, for every $y,Z\in\mathbb{k}^{*}$ satisfying $y\notin\{Z^{n}\mid n\in\mathbb{Z}\}$, there exists an augmentation $\varepsilon$ of the knot DGA of $K_{0}$ to $\mathbb{k}$ such that $\varepsilon(\mu)=y$ and $\varepsilon(U)=Z$.
There are two new points in algebraic arguments in this paper. Let us explain them by focusing on the knot $K^{\prime}\#T_{(2,2m+1)}$ for $m\in\mathbb{Z}_{\geq 1}$.
The first point is that we find a constraint on augmentations of some knot DGAs in terms of roots of polynomials in $\mathbb{k}[T]$, where $\mathbb{k}$ is any field. Actually, we find a specific polynomial $P_{m}\in\mathbb{Z}[\mu,U][T]$ for every $m\in\mathbb{Z}_{\geq 1}$, whose degree in $T$ is $m+1$, with the following property (see Proposition 3.10): For any augmentation $\varepsilon$ of the knot DGA of $K^{\prime}\#T_{(2,2m+1)}$ to $\mathbb{k}$ such that $\varepsilon(\mu)=y$ and $\varepsilon(U)=Z$, the polynomial $\left.P_{m}\right|_{\mu=y,U=Z}\in\mathbb{k}[T]$ has a root in $\mathbb{k}$. About the derivation of $P_{m}$, see Remark 1.3 below. We will focus on the case $\mathbb{k}=\mathbb{Q}$. It is worth noting that this constraint on augmentations is trivial if $\mathbb{k}$ is algebraically closed.
The second point is that we use a classical result in number theory. Indeed, Theorem 1.1 when $K=K^{\prime}\#T_{(2,2m+1)}$ for $m\in\mathbb{Z}_{\geq 1}$ is proved by a contradiction as follows: Suppose that $K_{0}=K^{\prime}\#T_{(2,2m+1)}$ and $K_{1}$ is the unknot. Then, Proposition 1.2 and the property of $P_{m}$ for $\mathbb{k}=\mathbb{Q}$ imply that $\left.P_{m}\right|_{\mu=2,U=Z}\in\mathbb{Q}[T]$ has a root in $\mathbb{Q}$ for every $Z\in\mathbb{Q}^{*}\setminus\{2^{\pm 1}\}$. On the other hand, by using Hilbert’s irreducibility theorem, we can show that there are infinitely many $Z\in\mathbb{Q}$ such that $\left.P_{m}\right|_{\mu=2,U=Z}\in\mathbb{Q}[T]$ does not have a root in $\mathbb{Q}$, which is a contradiction. (From the case of $m\geq 1$, we can deduce Theorem 1.1 when $K=K^{\prime}\#T_{(2,-2m-1)}$.)
The proof of Theorem 1.1 when $K=K^{\prime}\#4_{1}$ goes in a similar way. We explicitly give a polynomial $P\in\mathbb{Z}[\mu^{\pm},U^{\pm}][T]$, whose degree in $T$ is $3$, with the following property (see Proposition 3.17): For any augmentation $\varepsilon$ of the knot DGA of $K^{\prime}\#4_{1}$ to $\mathbb{k}$ such that $\varepsilon(\mu)=y$ and $\varepsilon(U)=Z$, the polynomial $\left.P\right|_{\mu=y,U=Z}\in\mathbb{k}[T]$ has a root in $\mathbb{k}$. Using concrete calculations and the rational root theorem, we show that $\left.P\right|_{\mu=-1,U=-2}\in\mathbb{Q}[T]$ does not have a root in $\mathbb{Q}$. Combining with Proposition 1.2 for $\mathbb{k}=\mathbb{Q}$, Theorem 1.1 when $K=K^{\prime}\#4_{1}$ is deduced.
@@
The polynomials $P_{m}$ for $K^{\prime}\#T_{(2,2m+1)}$ and $P$ for $K^{\prime}\#4_{1}$ are found from concrete computations on some part of (the homology of) the knot DGA which comes from the component $T_{(2,2m+1)}$ or $4_{1}$ of the connected sum. Our computation relies on braid representations of $T_{(2,2m+1)}$ and $4_{1}$, which are rather simple. It might be possible to find a polynomial with a similar property as $P_{m}$ for more complicated knots, such as the $(p,q)$-torus knot when $3\leq p<|q|$.
We also provide a theorem which is a refinement of [16, Theorem 1.1]. For an oriented knot $K$ in $\mathbb{R}^{3}$, let $V_{K}$ denote the augmentation variety defined by [15, Definition 5.1] as a subset of $(\mathbb{C}^{*})^{3}$. Let $(x,y,Z)$ be the coordinate of $\mathbb{C}^{3}$.
Theorem 1.4 (Corollary 3.7).
For oriented knots $K_{0}$ and $K_{1}$ in $\mathbb{R}^{3}$, suppose that there exists $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ intersects $\mathbb{R}^{3}$ cleanly along $K_{1}$. Then, there exist $a\in\{\pm 1\}$ and integers $b,c$ and $d$ for which we have an injective map
$V_{K_{1}}\hookrightarrow V_{K_{0}}\colon(x,y,Z)\mapsto(x^{a}y^{b}Z^{c},y^{a}Z^{d},Z).$
Remark 1.5.
@@
Identifying $\{(x,y,Z)\in(\mathbb{C}^{*})^{3}\mid Z=1\}$ with $(\mathbb{C}^{*})^{2}$, we obtain $V_{K}\cap\{Z=1\}\subset(\mathbb{C}^{*})^{2}$, which is a version of augmentation variety used in [16]. This corresponds to the reduction of the coefficient ring of the knot DGAs from $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$ to $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$ by setting $U=1$. It follows from Theorem 1.4 that there exists an injective map
$V_{K_{1}}\cap\{Z=1\}\hookrightarrow V_{K_{0}}\cap\{Z=1\}\colon(x,y)\mapsto(x^{a}y^{b},y^{a})$
@@
for some $a\in\{\pm 1\}$ and $b\in\mathbb{Z}$. This is a slight refinement of [16, Theorem 1.1], but it gives no constraint on $K_{0}$ if we suppose that $K_{1}$ is the unknot. The reason is that $V_{K_{1}}\cap\{Z=1\}=\{x=1\}\cup\{y=1\}$ if $K_{1}$ is the unknot (see [15, Example 3.13]) and $\{x=1\}\cup\{y=1\}$ is a subset of $V_{K_{0}}\cap\{Z=1\}$ for every knot $K_{0}$ (see [13, Proposition 5.6] and [15, Exercise 3.17]).
Notations. In this paper, for any manifold $X$, we take $\mathbb{Z}$ as the coefficient ring of the singular homology group of $X$ and denote it by $H_{*}(X)$, unless otherwise specified. Likewise, the relative homology group over $\mathbb{Z}$ is denoted by $H_{*}(X,A)$ for any submanifold $A$ of $X$.
@@
Organization of paper. In Section 2, for a contact manifold $Y$ and its Legendrian submanifold $\Lambda$ satisfying certain conditions, we introduce the Chekanov-Eliashberg DGA of $(Y,\Lambda)$ with coefficients in $\mathbb{Z}[H_{2}(Y,\Lambda)]$. We then define a DGA map associated to an exact Lagrangian cobordism in $\mathbb{R}\times Y$. In Section 2.5, we will see how to apply this DGA map to a setting with clean Lagrangian intersection as in Question 1.1. In Section 3, we combine the argument in Section 2.5 with the results of [3, 15] about the knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$. Proposition 1.2 and Theorem 1.4 are proved in Section 3.1. Theorem 1.1 for $K=K^{\prime}\#T_{(2,2m+1)}$ with $m\neq 0,-1$ (resp. $K=K^{\prime}\#4_{1}$) is proved by examining the knot DGA of $K^{\prime}\#T_{(2,2m+1)}$ in Section 3.2 (resp. $K^{\prime}\#4_{1}$ in Section 3.3). In Appendix A, by using Lagrangian Floer cohomology group over $\mathbb{F}_{2}[\lambda^{\pm}]$, we prove Proposition 2.7 which gives a homological constraint on $1$-dimensional clean intersections. This strengthen a result in [16]. Nevertheless, the present proof is much simpler than that of [16].

### v3

@@ Abstract
Let $K$ be a knot in $\mathbb{R}^{3}$ which has the $(2,q)$-torus knot for $q\neq\pm 1$ or the figure-eight knot as a component of connected sum. For its conormal bundle $L_{K}$ in $T^{*}\mathbb{R}^{3}$, we show that there is no compactly supported Hamiltonian diffeomorphism $\varphi$ on $T^{*}\mathbb{R}^{3}$ such that $\varphi(L_{K})$ intersects the zero section $\mathbb{R}^{3}$ cleanly along the unknot in $\mathbb{R}^{3}$. Using symplectic field theory, the proof is reduced to studying the augmentation variety $V_{\mathbb{k}}(K)$ of $K$ over a filed $\mathbb{k}$. The key point of this paper is finding an algebraic constraint on $V_{\mathbb{k}}(K)$ which is valid only when $\mathbb{k}$ is not algebraically closed, and the proof is completed by some arithmetic argument with $\mathbb{k}=\mathbb{Q}$.
@@ 1 Introduction
Backgraound. Let us consider the cotangent bundle $T^{*}\mathbb{R}^{3}$ of $\mathbb{R}^{3}$ with a canonical symplectic structure. We identify the zero section with $\mathbb{R}^{3}$. Given any knot $K$ in $\mathbb{R}^{3}$, let $L_{K}$ denote its conormal bundle. It is a Lagrangian submanifold of $T^{*}\mathbb{R}^{3}$ and intersects $\mathbb{R}^{3}$ cleanly along the knot $K$. Let $\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$ denote the set of Hamiltonian diffeomorphisms on $T^{*}\mathbb{R}^{3}$ associated to compactly supported Hamiltonians on $T^{*}\mathbb{R}^{3}$. Using these notations, we ask the following question, which was observed by the author in [15]: Is there $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K})$ intersects $\mathbb{R}^{3}$ cleanly along a knot whose isotopy class in $\mathbb{R}^{3}$ is different from $K$?
As a special case, the answer to this question when $K$ is the unknot is ‘no’. This can be deduced from the result by Ganatra-Pomerleano [8, Proposition 6.29] about clean Lagrangian intersections in a plumbing of two copies of the cotangent bundle of $S^{3}$ along the unknots (see [15, Section 1]). Smith-Wemyss [19, Proposition 4.10] and Asplund-Li [1, Theorem 1.3] also studied clean Lagrangian intersections in plumbings of two copies of $T^{*}S^{3}$ along the unknots. These results are concerned with Smith’s question on the ‘persistence (or rigidity) of unknottedness of Lagrangian intersections’ [8, Question 1.5].
In a general case, under the assumption that $\varphi(L_{K})$ for $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ intersects $\mathbb{R}^{3}$ cleanly along a knot $K^{\prime}$ in $\mathbb{R}^{3}$, the author proved a theorem [15, Theorem 1.1], which relates the augmentation variety in $(\mathbb{C}^{*})^{2}$ of $K$ with that of $K^{\prime}$. This result leads us to the same answer to the question as above when $K$ is the unknot [15, Theorem 1.3]. We can also deduce some constraints on $K^{\prime}$ when $K$ has certain non-trivial knot types [15, Theorem 1.4, 1.5].
In fact, the aforementioned result [15, Theorem 1.1] is useless to solve the following question which asks the persistence of knottedness of clean intersections. For the reason, see Remark 1.3.
Question 1.1.
@@
Main results. The goal of this paper is to give a negative answer to Question 1.1 when $K$ has a certain knot type by using augmentation variety over $\mathbb{Q}$ and some arithmetic argument. Let $T_{(2,2m+1)}$ for $m\in\mathbb{Z}$ denote the $(2,2m+1)$-torus knot and $4_{1}$ denote the figure-eight knot.
@@
For the proof, we use Theorem 1.2 below. For an oriented knot $K$ in $\mathbb{R}^{3}$, the knot DGA is defined combinatorially from a braid representation of a knot. In this paper, we refer to a definition in [13, 3, 14] with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$. Fix a field $\mathbb{k}$. From the set of augmentations of the knot DGA of $K$ to $\mathbb{k}$, following [14, Definition 5.1], we will define the augmentation variety $V_{\mathbb{k}}(K)\subset(\mathbb{k}^{*})^{3}$ in Definition 3.4. Let $(x,y,Z)$ be the coordinate of $\mathbb{k}^{3}$.
Theorem 1.2 (Corollary 3.5).
For $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ and oriented knots $K_{0}$ and $K_{1}$ in $\mathbb{R}^{3}$, suppose that $\varphi(L_{K_{0}})$ intersects $\mathbb{R}^{3}$ cleanly along $K_{1}$. Then, there exist $a\in\{\pm 1\}$ and integers $b,c$ and $d$ for which we have an injective map
$V_{\mathbb{k}}(K_{1})\hookrightarrow V_{\mathbb{k}}(K_{0})\colon(x,y,Z)\mapsto(x^{a}y^{b}Z^{c},y^{a}Z^{d},Z).$
This is a refinement of [15, Theorem 1.1], and the proof is parallel to [15] except an enhancement of the coefficient ring. We use the results by Ekholm-Etnyre-Ng-Sullivan [6, 3] which show that the Chekanov-Eliashberg DGA of the unit conormal bundle of $K_{i}$ is stable tame isomorphic to the knot DGA of $K_{i}$ for $i\in\{0,1\}$, and we compose the stable tame isomorphisms with a DGA map associated to an exact Lagrangian cobordism. For more on the coefficient ring, see Subsection 2.5.1.
@@
Identifying $\{(x,y,Z)\in(\mathbb{k}^{*})^{3}\mid Z=1\}$ with $(\mathbb{k}^{*})^{2}$, we obtain $V_{\mathbb{k}}(K)\cap\{Z=1\}\subset(\mathbb{k}^{*})^{2}$, which is a version of augmentation variety used in [15]. This corresponds to the reduction of the coefficient ring of the knot DGAs from $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$ to $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$ by setting $U=1$. It follows from Theorem 1.2 that there exists an injective map
$V_{\mathbb{k}}(K_{1})\cap\{Z=1\}\hookrightarrow V_{\mathbb{k}}(K_{0})\cap\{Z=1\}\colon(x,y)\mapsto(x^{a}y^{b},y^{a})$
@@
for some $a\in\{\pm 1\}$ and $b\in\mathbb{Z}$. This is a slight refinement of [15, Theorem 1.1], but it gives no constraint on $K_{0}$ if $K_{1}=O$, where $O$ is the unknot. The reason is that a relation $V_{\mathbb{k}}(O)\cap\{Z=1\}\subset V_{\mathbb{k}}(K)\cap\{Z=1\}$ automatically holds for any knot $K$. See [12, Section 5.2]
Compared with [15], where $V_{\mathbb{C}}(K_{i})\cap\{Z=1\}\subset(\mathbb{C}^{*})^{2}$ for $i\in\{0,1\}$ were used, a novel point of this paper lies in an algebraic argument to deduce Theorem 1.1 from Theorem 1.2.
Let us explain the idea in the case $K=K^{\prime}\#T_{(2,2m+1)}$ for $m\in\mathbb{Z}_{\geq 1}$. It suffices to show that for some field $\mathbb{k}$, $V_{\mathbb{k}}(O)$ cannot be embedded into $V_{\mathbb{k}}(K^{\prime}\#T_{(2,2m+1)})$ by the map of Theorem 1.2. From the knot DGA of $O$, which is quite simple, it is easy to compute that $V_{\mathbb{k}}(O)=\{Z-x-y+xy=0\}$. However, determining $V_{\mathbb{k}}(K^{\prime}\#T_{(2,2m+1)})$ is hard because the knot DGA is complicated in general. As a matter of fact, for the closure of a braid with $n$ strands, the 0-th homology of its knot DGA has n(n1)n(n-1) generators and $2n^{2}$ relations. See Subsection 3.2.1.
To solve this problem, we use an algebraic constraint on $V_{\mathbb{k}}(K^{\prime}\#T_{(2,2m+1)})$. In Section 3.2, by focusing on specific two relations in the 0-th homology, we will extract a polynomial $P_{m}\in\mathbb{Z}[\mu^{\pm},U^{\pm}][T]$ defined in an explicit way with the following property:
$V_{\mathbb{k}}(K^{\prime}\#T_{(2,2m+1)})\subset\{(x,y,Z)\mid\left.P_{m}\right|_{\mu=y,U=Z}\in\mathbb{k}[T]\text{ has a root in }\mathbb{k}\}.$
Then, the proof of Theorem 1.1 is reduced to finding $(x,y,Z)\in V_{\mathbb{k}}(O)$ such that the $\mathbb{k}$-coefficient polynomial $\left.P_{m}\right|_{\mu=y^{a}Z^{d},U=Z}$ does not have a root in $\mathbb{k}$. It is worth noting that this approach does not work if $\mathbb{k}$ is algebraically closed. Actually, we take $\mathbb{k}=\mathbb{Q}$ and apply Hilbert’s irreducibility theorem in the proof of Theorem 3.14.
In the case $K=K^{\prime}\#4_{1}$, we can explicitly find a polynomial (of degree $3$ in $T$) with a similar property. We take $\mathbb{k}=\mathbb{Q}$ and use the rational root theorem in the proof of Theorem 3.17.
@@
Organization of paper. In Section 2, for a contact manifold $Y$ and its Legendrian submanifold $\Lambda$ satisfying certain conditions, we introduce the Chekanov-Eliashberg DGA of $(Y,\Lambda)$ with coefficients in $\mathbb{Z}[H_{2}(Y,\Lambda)]$. We then define a DGA map associated to an exact Lagrangian cobordism in $\mathbb{R}\times Y$. In Section 2.5, we will see how to apply this DGA map to a setting with clean Lagrangian intersection. In Section 3, we combine the argument in Section 2.5 with the results of [3, 14] about the knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$. Theorem 1.2 is proved in Section 3.1. Theorem 1.1 for $K=K^{\prime}\#T_{(2,2m+1)}$ with $m\neq 0,-1$ (resp. $K=K^{\prime}\#4_{1}$) is proved by examining the knot DGA of $K^{\prime}\#T_{(2,2m+1)}$ in Section 3.2 (resp. $K^{\prime}\#4_{1}$ in Section 3.3). In Appendix A, by using Lagrangian Floer cohomology group over $\mathbb{F}_{2}[\lambda^{\pm}]$, we prove Proposition 2.7 which gives a homological constraint on $1$-dimensional clean intersections. This strengthen a result in [15, Section 2]. Nevertheless, the present proof is much simpler than that of [15].
Notations. In this paper, we take $\mathbb{Z}$ as the coefficient ring of the singular homology groups, unless otherwise mentioned.